# appAppendix References (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v2

@@ Abstract
The diffusion model has long been plagued by scalability and quadratic complexity issues, especially within transformer-based structures. In this study, we aim to leverage the long sequence modeling capability of a State-Space Model called Mamba to extend its applicability to visual data generation. Firstly, we identify a critical oversight in most current Mamba-based vision methods, namely the lack of consideration for spatial continuity in the scan scheme of Mamba. Secondly, building upon this insight, we introduce Zigzag Mamba, a simple, plug-and-play, minimal-parameter burden, DiT style solution, which outperforms Mamba-based baselines and demonstrates improved speed and memory utilization compared to transformer-based baselines. Lastly, we integrate Zigzag Mamba with the Stochastic Interpolant framework to investigate the scalability of the model on large-resolution visual datasets, such as FacesHQ $1024\times 1024$ and UCF101, MultiModal-CelebA-HQ, and MS COCO $256\times 256$.
@@
Diffusion Model State-Space Model Stochastic Interpolant
@@ 1 Introduction
Diffusion models have demonstrated significant advancements across various applications, including image processing [67], video analysis [40], point cloud processing [87], and human pose estimation [29]. Many of these models are built upon Latent Diffusion Models (LDM)[67], which are typically based on the UNet backbone. However, scalability remains a significant challenge in LDMs[41]. Recently, transformer-based structures have gained popularity due to their scalability [65, 9] and effectiveness in multi-modal training [10]. Notably, the transformer-based structure DiT [65] has even contributed to enhancing the high-fidelity video generation model SORA [64] by OpenAI. Despite efforts to alleviate the quadratic complexity of the attention mechanism through techniques such as windowing [58], sliding [12], sparsification [18, 46], hashing [19, 74], Ring Attention [53, 14], Flash Attention [21] or a combination of them [8, 97], it remains a bottleneck for diffusion models.
On the other hand, State-Space Models [32, 35, 31] have shown great potential for long sequence modeling, competing with transformer-based methods. Several methods [30, 32, 27, 69] have been proposed to enhance the robustness [92], scalability [30], and efficiency [32, 33] of State-Space Models. Among these, a method called Mamba [30] aims to alleviate these issues through work-efficient parallel scanning and other data-dependent innovations. However, the advantage of Mamba lies in 1D sequence modeling, and extending it to 2D images is a challenging question. Previous works [96, 57] have proposed flattening 2D tokens directly by computer hierarchy such as row-and-column-major order, but this approach neglects Spatial Continuity, as shown in Figure 1. Other works [54, 60] consider various directions in a single Mamba block, but this introduces additional parameters and GPU memory burden. In this paper, we aim to emphasize the importance of Spatial Continuity in Mamba and propose several intuitive and simple methods to enable the application of Mamba blocks to 2D images by incorporating continuity-based inductive biases in images. We also generalize these methods to 3D with spatial-temporal factorization on 3D sequence.
In summary, our contributions are as follows: Firstly, we identify the critical issue of Spatial Continuity in generalizing the Mamba block from 1D sequence modeling to 2D image and 3D video modeling. Building on this insight, we propose a simple, plug-and-play, zero-parameter paradigm named Zigzag Mamba (ZigMa) that leverages spatial continuity to maximally incorporate the inductive bias from visual data. Secondly, we extend the methodology from 2D to 3D by factorizing the spatial and temporal sequences to optimize performance. Secondly, we provide comprehensive analysis surrounding the Mamba block within the regime of diffusion models. Lastly, we demonstrate that our designed Zigzag Mamba outperforms related Mamba-based baselines, representing the first exploration of Stochastic Interpolants on large-scale image data ($1024\times 1024$) and videos.
@@ 2 Related Works
Mamba. Several works [83, 82] have demonstrated that the State-Space Model possesses universal approximation ability under certain conditions. Mamba, as a new State-Space Model, has superior potential for modeling long sequences efficiently, which has been explored in various fields such as medical imaging [89, 60, 86, 68], image restoration [34, 95], graphs [11], NLP word byte [80], tabular data [2], point clouds [49], and image generation [25]. Among them, the most related to us are VisionMamba[96, 57], S4ND [63] and Mamba-ND [48]. VisionMamba[96, 57] uses a bidirectional SSM in discriminative tasks which incurs a high computational cost. Our method applies a simple alternative mamba diffusion in generative models. S4ND [63] introduces local convolution into Mamba’s reasoning process, moving beyond the use of only 1D data. Mamba-ND [48] takes multi-dimensionality into account in discriminative tasks, making use of various scans within a single block. In contrast, our focus is on distributing scan complexity across every layer of the network, thus maximizing the incorporation of inductive bias from visual data with zero parameter burden.
@@
Backbones in Diffusion Models. Diffusion models primarily employ UNet-based [39, 67] and ViT-based [65, 9] backbones. While UNet is known for high memory demands [67], ViT benefits from scalability [22, 17] and multi-modal learning [10]. However, ViT’s quadratic complexity limits visual token processing, prompting studies towards mitigating this issue [84, 21, 12]. Our work, inspired by Mamba [30], explores an SSM-based model as a generic diffusion backbone, retaining ViT’s modality-agnostic and sequential modeling advantages. Concurrently, DiffSSM [90] concentrates on unconditional and class conditioning within the S4 model [32]. DIS [25] mainly explores the state-space model on a relatively small scale, which is not the exact focus of our work. Our work significantly differs from theirs as it primarily focuses on the backbone design using the Mamba block and extends it to text conditioning. Furthermore, we apply our method to more complex visual data.
@@
SDE and ODE in Diffusion models. The realm of Score-based Generative Models encompasses significant contributions from foundational works such as Score Matching with Langevin Dynamics (SMLD) by Song et al.[71], and the advent of Diffusion Models with Denoising Score Matching (DDPMs) proposed by Ho et al.[39]. These methodologies operate within the framework of Stochastic Differential Equations (SDEs), a concept further refined in the research of Song et al. [72]. Recent research strides, as exemplified by Karras et al.[42] and Lee et al.[47], have showcased the efficacy of employing Ordinary Differential Equation (ODE) samplers for diffusion SDEs, offering significant reductions in sampling costs compared to traditional approaches that entail discretizing diffusion SDEs. Furthermore, within the domain of Flow Matching methodologies [51] and Rectified Flow architectures [55], both SMLD and DDPMs emerge as specialized instances under distinct paths of the Probability Flow ODE framework [72]. These models typically utilize velocity field parameterizations employing the linear interpolant, a concept that finds broader applications in the Stochastic Interpolant framework [3], with subsequent generalizations extending to manifold settings [13]. The SiT model [61] scrutinizes the interplay between interpolation methods in both sampling and training contexts, albeit in the context of smaller resolutions such as $512\times 512$. Our research endeavors to extend these insights to a larger scale, focusing on the generalization capabilities for 2D images of $1024\times 1024$ and 3D video data.
@@ 3.1 Background: State-Space Models
Recently, Mamba [30] largely improved the flexibility of SSMs by relaxing the time-invariance constraint on SSM parameters, while maintaining computational efficiency. By employing a work-efficient parallel scan, Mamba mitigates the impact of the sequential nature of recurrence, whereas fusing GPU operations removes the requirement to materialize the expanded state. In this paper, we focus on exploring the scanning scheme of Mamba in diffusion models to maximize the use of inductive-bias from multi-dimensional visual data.
@@ 3.2 Diffusion Backbone: Zigzag Mamba
DiT-Style Network. We opt to use the framework of ViT by AdaLN [65] rather than the skip-layer focused U-ViT structure [9], as ViT has been validated as a scalable structure in literature [64, 10, 17]. Considering the aforementioned points, it informs our Mamba network design depicted in Figure 4. The core component of this design is the Zigzag Scanning, which will be explained in the following paragraph.
@@
$\bar{{\mathbf{z}}}_{\Omega_{i}}$ $=\texttt{scan}({\mathbf{z}}_{\Omega}),$ (2)
@@
Figure 4: The Detail of our Zigzag Mamba block. The detail of Mamba Scan is shown in Figure 2. The condition can include a timestep and a text prompt. These are fed into an MLP, which separately modulates the Mamba scan for long sequence modeling and cross attention for multi-modal reasoning.
@@
(b) 3D Zigzag: Compared with we formulate the 2D zigzag in previous subsections, we follow the similar design to generalize it to 3D Zigzag to keep the continuity in 2D and 3D simultaneously. Potentially, the scheme has much more complexity. We heuristically list 8 schemes as well. However, we empirically find that this scheme will lead to suboptimal optimization.
@@ 4 Experiment
In this section, we begin by detailing the experimental setup concerning image and video datasets, as well as our training details. Subsequently, we delve into several in-depth analyses aimed at elucidating the rationale behind our method design across various resolutions. Finally, we present our results obtained from higher-resolution and more complex datasets.
@@ 4.1 Dataset and Training Detail
For text-conditioned generation, we conduct the experiments on the MultiModal-CelebA $256^{2}$,$512^{2}$ [88] and MS COCO $256\times 256$ [50] datasets. Both datasets are composed of text-image pairs for training. Typically, there are 5 to 10 captions per image in COCO and MultiModal-CelebA. We convert discrete texts to a sequence of embeddings using a CLIP text encoder [66] following Stable Diffusion [67]. Then these embeddings are fed into the network as a sequence of tokens.
Video dataset. UCF101 dataset is consists of 13,320 video clips, which are classified into 101 categories. The total length of these video clips is over 27 hours. All these videos are collected from YouTube and have a fixed frame rate of 25 FPS with the resolution of $320\times 240$. We randomly sample continuous 16 frames and resize the frames to $256\times 256$.
@@
MultiModal-CelebA256256256256 MultiModal-CelebA512512512512
@@ 4.2 Ablation Study
Spatial Continuity is Critical. We first explore the importance of spatial continuity in Mamba design by grouping patches of size $N\times N$ into various sizes: $2\times 2$, $4\times 4$, $8\times 8$, and $16\times 16$, resulting in groups of patch sizes $N/2\times N/2$, $N/4\times N/4$, $N/8\times N/8$, and $N/16\times N/16$, respectively. Then, we apply our designed Zigzag-8 scheme at the group level instead of the patch level. Figure 6 illustrates that with increased spatial continuity, notably improved performance is achieved. Furthermore, we compare our approach with random shuffling of $N\times N$ patches, revealing notably inferior performance under random shuffling conditions. All of these results collectively indicate that spatial continuity is a critical requirement when applying Mamba in 2D sequences.
Figure 6: Spatial Continuity Analysis. As we incrementally enlarge the patch group size, the continuous segment of the patch also expands. This enhances spatial continuity, which we find improves FID on MultiModal-CelebA 256, 512 dataset.
Furthermore, in Figure 7 (c) we conduct an ablation study on GPU memory and FPS among different variants of our method. We find that our method incurs nearly zero FPS and GPU memory burden when gradually increasing the Mamba Scan Schemes. This analysis provides insights into the comparative performance and efficiency of different diffusion backbones, highlighting the advantages of our proposed method.
Ablation study about the Order Receptive Field and FPS/GPU-Memory. As depicted in Fig. 10, Zigzag Mamba consistently maintains its GPU memory consumption and FPS rate, even with a gradually increasing Order Receptive Field. In contrast, our primary baseline, Parallel Mamba, along with variants like Bidirectional Mamba and Vision Mamba [57, 96], experience a consistent decrease in FPS due to increased parameters. Notably, Zigzag Mamba, with an Order Receptive Field of 8, can perform faster without altering parameters.
@@
(c) GPU usage of variants of our method.
(d) FPS v.s. Patch Size.
Figure 7: (a, b).GPU Memory usage and FPS between our method and transformer-based methods(U-VIT [9] and DiT [65]). (c). GPU memory usage and FPS among various variants of our method. (d). The relationship between patch size and FID under various resolutions.
(a) Order Receptive Field v.s. FPS.
(b) Order Receptive Field v.s. GPU Memory.
Figure 8: The ablation study about Order Receptive Field, FPS, GPU Memory.
Patch size. We conducted an ablation study on patch sizes ranging from 1, 2, 4, to 8 in Figure 7 (d), aiming to explore their behaviors under the framework of Mamba. The results reveal that the FID deteriorates as the patch size increases, aligning with the common understanding observed in the field of transformers [23, 78]. This suggests that smaller patch sizes are crucial for optimal performance.
Table 2: Main Results on MS-COCO dataset. Our method consistently outperforms the baseline and can achieve even better results when the training scale is increased.
@@
Zigzag-8 $\times$ 16GPU 33.8
Table 3: Main result on FacesHQ-1024 dataset with 4,094 tokens in latent space. Our method can outperform the baseline and can achieve even better results when the training scale is increased.
@@
Bidirection Mamba-16GPU [96] 51.1 66.3
Zigzag-Mamba -16GPU 37.8 50.5
Zigzag-Mamba -32GPU 26.6 31.2

### v3

@@ Abstract
The diffusion model has long been plagued by scalability and quadratic complexity issues, especially within transformer-based structures. In this study, we aim to leverage the long sequence modeling capability of a State-Space Model called Mamba to extend its applicability to visual data generation. Firstly, we identify a critical oversight in most current Mamba-based vision methods, namely the lack of consideration for spatial continuity in the scan scheme of Mamba. Secondly, building upon this insight, we introduce Zigzag Mamba, a simple, plug-and-play, minimal-parameter burden, DiT style solution, which outperforms Mamba-based baselines and demonstrates improved speed and memory utilization compared to transformer-based baselines, also this heterogeneous layerwise scan enables zero memory and speed burden when we consider more scan paths. Lastly, we integrate Zigzag Mamba with the Stochastic Interpolant framework to investigate the scalability of the model on large-resolution visual datasets, such as FacesHQ $1024\times 1024$ and UCF101, MultiModal-CelebA-HQ, and MS COCO $256\times 256$.
@@
Diffusion Model State-Space Model Stochastic Interpolants
@@ 1 Introduction
Diffusion models have demonstrated significant advancements across various applications, including image processing [48, 45, 84], video analysis [44], point cloud processing [109], representation learning [30] and human pose estimation [32]. Many of these models are built upon Latent Diffusion Models (LDM)[84], which are typically based on the UNet backbone. However, scalability remains a significant challenge in LDMs[50]. Recently, transformer-based structures have gained popularity due to their scalability [80, 9] and effectiveness in multi-modal training [10]. Notably, the transformer-based structure DiT [80] has even contributed to enhancing the high-fidelity video generation model SORA [78] by OpenAI. Despite efforts to alleviate the quadratic complexity of the attention mechanism through techniques such as windowing [71], sliding [13], sparsification [19, 56], hashing [20, 93], Ring Attention [66, 15], Flash Attention [23] or a combination of them [8, 124], it remains a bottleneck for diffusion models.
On the other hand, State-Space Models [35, 39, 34] have demonstrated significant potential for long sequence modeling, rivaling transformer-based methods. Their biological similarity [95] and efficient memory state also advocate for the use of the State-Space model over the transformer. Several methods [33, 35, 29, 88] have been proposed to enhance the robustness [116], scalability [33], and efficiency [35, 36] of State-Space Models. Among these, a method called Mamba [33] aims to alleviate these issues through work-efficient parallel scanning and other data-dependent innovations. However, the advantage of Mamba lies in 1D sequence modeling, and extending it to 2D images is a challenging question. Previous works [123, 70] have proposed flattening 2D tokens directly by computer hierarchy such as row-and-column-major order, but this approach neglects Spatial Continuity, as shown in Figure 1. Other works [67, 73] consider various directions in a single Mamba block, but this introduces additional parameters and GPU memory burden. In this paper, we aim to emphasize the importance of Spatial Continuity in Mamba and propose several intuitive and simple methods to enable the application of Mamba blocks to 2D images by incorporating continuity-based inductive biases in images. We also generalize these methods to 3D with spatial-temporal factorization on 3D sequence.
In summary, our contributions are as follows: Firstly, we identify the critical issue of Spatial Continuity in generalizing the Mamba block from 1D sequence modeling to 2D image and 3D video modeling. Building on this insight, we propose a simple, plug-and-play, zero-parameter heterogeneous layerwise scan paradigm named Zigzag Mamba (ZigMa) that leverages spatial continuity to maximally incorporate the inductive bias from visual data. Secondly, we extend the methodology from 2D to 3D by factorizing the spatial and temporal sequences to optimize performance. Secondly, we provide comprehensive analysis surrounding the Mamba block within the regime of diffusion models. Lastly, we demonstrate that our designed Zigzag Mamba outperforms related Mamba-based baselines, representing the first exploration of Stochastic Interpolants on large-scale image data ($1024\times 1024$) and videos.
@@ 2 Related Works
Mamba. Several works [103, 102, 103] have demonstrated that the State-Space Model possesses universal approximation ability under certain conditions. Mamba, as a new State-Space Model, has superior potential for modeling long sequences efficiently, which has been explored in various fields such as medical imaging [111, 73, 108, 86], video [58, 79], image restoration [38, 122], graphs [12], NLP word byte [100], tabular data [2], point clouds [61], human motion [120, 106], multi-task [62] and image generation [27]. Among them, the most related to us are VisionMamba[123, 70], S4ND [77] and Mamba-ND [59]. VisionMamba[123, 70] uses a bidirectional SSM in discriminative tasks which incurs a high computational cost. Our method applies a simple alternative mamba diffusion in generative models. S4ND [77] introduces local convolution into Mamba’s reasoning process, moving beyond the use of only 1D data. Mamba-ND [59] takes multi-dimensionality into account in discriminative tasks, making use of various scans within a single block. In contrast, our focus is on distributing scan complexity across every layer of the network, thus maximizing the incorporation of inductive bias from visual data with zero parameter burden. Scan curve is an important direction in SSM, PointMamba [61] is a representative work that employs SSM with space curves (e.g., Hilbert) for point cloud analysis, achieving remarkable performance. In contrast with them, our preliminary results show that the Hilbert curve doesn’t work well with our method (see Appendix), while our method can be regarded as the simplest Peano curve. For more information related to Mamba’s work, please refer to the survey [105].
@@
Backbones in Diffusion Models. Diffusion models primarily employ UNet-based [43, 84] and ViT-based [80, 9] backbones. While UNet is known for high memory demands [84], ViT benefits from scalability [24, 18] and multi-modal learning [10]. However, ViT’s quadratic complexity limits visual token processing, prompting studies towards mitigating this issue [104, 23, 13]. Our work, inspired by Mamba [33], explores an SSM-based model as a generic diffusion backbone, retaining ViT’s modality-agnostic and sequential modeling advantages. Concurrently, DiffSSM [112] concentrates on unconditional and class conditioning within the S4 model [35]. DIS [27] mainly explores the state-space model on a relatively small resolution, which is not the exact focus of our work. Our work significantly differs from theirs as it primarily focuses on the backbone design using the Mamba block and extends it to text conditioning. Furthermore, we apply our method to more complex visual data.
@@
SDE and ODE in Diffusion models. The realm of Score-based Generative Models encompasses significant contributions from foundational works such as Score Matching with Langevin Dynamics (SMLD) by Song et al.[90], and the advent of Diffusion Models with Denoising Score Matching (DDPMs) proposed by Ho et al.[43]. These methodologies operate within the framework of Stochastic Differential Equations (SDEs), a concept further refined in the research of Song et al. [91]. Recent research strides, as exemplified by Karras et al.[52] and Lee et al.[57], have showcased the efficacy of employing Ordinary Differential Equation (ODE) samplers for diffusion SDEs, offering significant reductions in sampling costs compared to traditional approaches that entail discretizing diffusion SDEs. Furthermore, within the domain of Flow Matching [64] and Rectified Flow [68], both SMLD and DDPMs emerge as specialized instances under distinct paths of the Probability Flow ODE framework [91], with broad applications in vision [49, 28, 22], depth [37], human motion [47], even language [46]. These models typically utilize velocity field parameterizations employing the linear interpolant, a concept that finds broader applications in the Stochastic Interpolant framework [3], with subsequent generalizations extending to manifold settings [14]. The SiT model [74] scrutinizes the interplay between interpolation methods in both sampling and training contexts, albeit in the context of smaller resolutions such as $512\times 512$. Our research endeavors to extend these insights to a larger scale, focusing on the generalization capabilities for 2D images of $1024\times 1024$ and 3D video data.
@@ 3.1 Background: State-Space Models
Recently, Mamba [33] largely improved the flexibility of SSMs in Language Modelling by relaxing the time-invariance constraint on SSM parameters, while maintaining computational efficiency. Several studies [123, 70] have been conducted to adapt the use of Mamba from unidimensional language data to multidimensional visual data. While most of these studies try to duplicate the A to facilitate the new (reversed) direction, this approach can lead to additional parameters and an increased memory burden. In this paper, we focus on exploring the scanning scheme of Mamba in diffusion models to efficiently maximize the use of inductive-bias from multi-dimensional visual data with zero parameter and memory burden.
@@ 3.2 Diffusion Backbone: Zigzag Mamba
DiT-Style Network. We opt to use the framework of DiT by AdaLN [80] rather than the skip-layer focused U-ViT structure [9], as DiT has been validated as a scalable structure in literature [78, 10, 18]. Additionally, the Hourglass structure with downsampling [76, 85] requires selecting the depth and width based on the complexity of the dataset and task. This requirement limits the flexibility of the solution. Considering the aforementioned points, it informs our Mamba network design depicted in Figure 4. The core component of this design is the Zigzag Scanning, which will be explained in the following paragraph.
@@
$\bar{{\mathbf{z}}}_{\Omega_{i}}$ $=\texttt{scan}({\mathbf{z}}_{\Omega_{i}}),$ (2)
@@
Figure 4: The Detail of our Zigzag Mamba block. The detail of Mamba Scan is shown in Figure 2. The condition can include a timestep and a text prompt. These are fed into an MLP, which separately modulates the Mamba scan for long sequence modeling and cross-attention for multi-modal reasoning.
@@
(b) 3D Zigzag: Compared with the formulation of the 2D zigzag in previous subsections, we follow the similar design to generalize it to 3D Zigzag to keep the continuity in 2D and 3D simultaneously. Potentially, the scheme has much more complexity. We heuristically list 8 schemes as well. However, we empirically find that this scheme will lead to suboptimal optimization.
@@ 4.1 Dataset and Training Detail
MultiModal-CelebA-256 MultiModal-CelebA-512
@@
Video Dataset. UCF101 dataset consists of 13,320 video clips, which are classified into 101 categories. The total length of these video clips is over 27 hours. All these videos are collected from YouTube and have a fixed frame rate of 25 FPS with the resolution of $320\times 240$. We randomly sample continuous 16 frames and resize the frames to $256\times 256$.
@@ 4.2 Ablation Study
Table 2: Ablation about Position Embedding (PE) on unconditional CelebA dataset ($256^{2}$). To better abate PE and eliminate the conditional signal’s influence, we use an unconditional dataset.
FID/FDD $\downarrow$ No PE Cosine PE Learnable PE
VisionMamba [123] 21.33/21.00 18.47/19.90 16.38/18.20
ZigMa 14.27/18.00 14.04/17.91 13.32/17.40
@@
(c) Order Receptive Field v.s. GPU Memory.
(d) Order Receptive Field v.s. FPS.
Figure 6: (a, b).GPU Memory usage and FPS between our method and transformer-based methods(U-VIT [9] and DiT [80]). (c). Order Receptive Field and GPU memory (d). Order Receptive Field and FPS. Order Receptive Field denotes how many scan paths we consider in our network design.
Ablation about Position Embedding. As shown in Table 2, the learnable embedding performs better than the Sinusoidal embedding, which in turn performs better than no position embedding. In various cases, our zigzag method surpasses the baselines. Notably, our performance remains almost unchanged whether we use the Sinusoidal position embedding or no position embedding. This suggests that our method can better incorporate spatial inductive-bias compared to our baseline. Finally, using the learnable position embedding provides further, albeit marginal, gains suggesting that better position embedding exists even within our zigzag scan scheme. We find that  [79] shares the same conclusion as us in video-related tasks.
Table 3: Main result on FacesHQ-1024 dataset with 4,094 tokens in latent space and bs=512. Our method can outperform the baseline and can achieve even better results when the training scale is increased.
[…]
VisionMamba [123] 51.1 66.3
ZigMa 37.8 50.5
ZigMa bs $\times$ 2 26.6 31.2
Table 4: Main Results on MS-COCO dataset with bs=256. Our method consistently outperforms the baseline. ZigMa with 8 scans performs much better compared with the baseline.
@@
Table 5: Transformer-based methods comparison on unconditional CelebA256.
@@
U-ViT 14.50 35.10 12.5
DiT 14.64 29.20 5.5
ZigMa 14.27 17.80 5.2
Table 6: Video Scan Scheme on UCF101 dataset with bs=32.
[…]
Bidirection [123] 256.1 320.2
3D Zigzag 238.1 282.3
Our 216.1 210.2
Bidirection [123] bs×$\times$4 146.2 201.1
ZigMa bs×$\times$4 121.2 140.1
Ablation study about the Order Receptive Field and FPS/GPU-Memory. As depicted in Fig. 6 (c,d), Zigzag Mamba consistently maintains its GPU memory consumption and FPS rate, even with a gradually increasing Order Receptive Field. In contrast, our primary baseline, Parallel Mamba, along with variants like Bidirectional Mamba and Vision Mamba [70, 123], experience a consistent decrease in FPS due to increased parameters. Notably, Zigzag Mamba, with an Order Receptive Field of 8, can perform faster without altering parameters.
Comparison with transformer-based methods. We show the result in Table 6 on unconditional generation task.Our method achieves performance comparable to Transformer-based methods, with significantly less memory consumption and fewer FLOPS.